# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v2

@@ VI Perspective on wormhole evaporation over long times
As shown in the previous section, charge and phase fluctuations accumulate very rapidly due to the fundamental instability inherent to the wormhole horizons or interior (in case of nearest neighbour inductive coupling). In particular, the more the system accumulates charge and phase fluctuations over time, the more the quadratic approximation of the Josephson energy becomes inaccurate, such that predicting the system dynamics for long times becomes a much harder task (as the system can no longer be approximated by non-interacting bosons). This is well beyond the scope of the present work. We nonetheless find it illustrative to speculate on a qualitative level about the long term fate of the wormhole – especially as it allows us to distinguish between intrinsic (spontaneous) collapse, and dissipative relaxation. In addition, some of the resulting concepts provide an interesting segue to the subsequent idea: quantum superpositions of the spacetime geometry.
@@
To this end, let us take into account interactions with an environment. For the sake of simplicity and concreteness, we consider the phase difference $\phi_{j}-\phi_{j-1}+\phi_{\text{ext},j}$ across a single junction. At the beginning of the wormhole quench, this phase difference is either localized around 00 (minimum of the cosine, if the junction is positioned outside the wormhole) or around ±$\pm\pi$ (maximum of the cosine, inside the wormhole). In the presence of the instability, the quantum fluctuations of this phase difference blow up over time, as shown in the previous section. Now suppose that the environment entangles (either weakly or strongly) with the current across this junction. This means that the phase difference gets spontaneously projected onto a more localized state. Notably, this type of environment-induced process thus diminishes quantum fluctuations. We note that the same picture holds in perfect analogy for the accumulation of charge noise quantum fluctuations, respectively, the reduction thereof by a dissipative process extracting information about the charge.
@@
To proceed, note that the location of the wave function now is no longer at 00 or $\pi$, but can be (with a finite probability) at a certain distance from the minimum or maximum. With the new location updated, we again quadratically expand the cosine around the new peak position of the phase difference. Importantly, due to the cosine behaviour of the junction the resulting effective inductance is now different (the curvature of the cosine obviously changes as a function of the position at which it is calculated). Thus, as time progresses, and the phase difference starts to classically diffuse after repeated environment-induced measurements, the effective inductance at a given junction likewise fluctuates classically. At this point, the impact of such dissipative processes could likewise be detectable in time-of-flight measurements of wave packets within the array. At any rate, the most likely long-term outcome is that all phase differences relax to the ground state, equivalent to the complete evaporation of the analog wormhole. We note that for a system on a ring (with periodic boundaries), or a system with a finite mass term (Cooper-pair leakage to ground), the emergence of soliton states might be a possibility.
@@
Let us repeat that even though the environment will undoubtedly have a big impact on the system dynamics (especially beyond the immediate transient time scale), the experimental signatures could not be more different. Analog Hawking radiation leads to an increase of quantum fluctuations, whereas the environment induces classical fluctuations. Thus, the observation of the former can distinguish between how the system intrinsically reacts to the creation of an instability, as compared to the impact of external perturbations.
@@
To conclude this section, we consider the very hypothetical case of negligible coupling to the environment, as it reveals an interesting additional idea which we pursue in the remainder of this work. If we assume that the build-up of quantum fluctuations of the phase could progress unimpeded over longer times, the nonlinearity of the Josephson energies results not in classical fluctuations of the effective inductance (as in the previous paragraph), but in a system that has to be interpreted as being in a quantum superposition of different effective inductances – a form of quantum superposition of the spacetime geometry itself. However, within the above considered setting, this effect is likely not to survive for very long due to environment-induced decoherence of the quantum fluctuations, and even for a highly protected system, it would be hard to analyse the effect in an unambiguous way. In what follows now, we provide a modification of this idea, not relying on instabilities (and thus not requiring any ultra-fast quenches), but instead on more general notions of Josephson junctions with multiple minimas. This endeavour will on the one hand introduce the notion of a quantum inductance, and with it, a likely more stable version of a quantum superposition of the analog spacetime.
VII Bistable Josephson effect and quantum spacetime geometry
As announced, in this section we propose a possible way to create and observe superpositions of two different spacetime geometries using quantum circuits. As already hinted at in the introduction, the idea that spacetime geometries can exist as a superposition within actual gravity is exciting, but at the same time highly contentious [54].
Note, moreover, that the currently most debated possibility is the observation of entanglement created between two quantum systems via the gravitational force, hence proving the quantum nature of gravitation [50, 51]. Since gravity, in part, is a theory of the structure of spacetime, it being quantum in nature may imply that spacetime can exist in quantum superposition. The notion that we introduce now is insofar related, as it allows for the formulation of a quantum superposition of analog spacetime geometry, by exploiting internal degrees of freedom of general non-linear inductor elements.
The central idea can be formulated as follows. Previously, the local signal speed in a circuit composed of Josephson junctions, is uniquely defined by the curvature of the minimum of the Josephson potential. However, if we use a potential with more than one minimum, where – crucially – each minimum has a different curvature, they will have different signal speed. Since these two minima can be considered a quantum double well, it is possible to prepare a state where one occupies a superposition of the two minima, ultimately allowing us to create a quantum superposition of signals travelling simultaneously at multiple different speeds.
[…]
Generalized Josephson junctions with a more complicated energy-phase relationship, with, e.g., multiple minima (so-called multistable Josephson junctions) have been studied for some time by now. The simplest one is a junction whose dominant process consists of the tunnelling of pairs of Cooper-pairs, described by an energy with the phase dependence $\sim\cos\left(2\phi\right)$ [100, 101]. Their design as proposed in Ref. [101] is shown in Fig. 8(a). This design can also be generalized to get an element with the phase dependence $\sim\cos\left(3\phi\right)$ (Fig. 8(b)) [123]. We now imagine a parallel shunt of a regular Josephson junction with these generalized junctions to realize a potential of the form (Fig. 7)
$V\sim-\cos 2\phi+\delta\cos\phi-\delta\cos 3\phi\,,$ (27)
with $|\delta|<1/2$. This function has two minima, whose curvatures are $\sim 4(1\pm 2\delta)$. We note that while the parallel shunt of the above three different junction types is very instructive to demonstrate the principle, the actual experimental fabrication of a potential of the form of Eq. (27) may be simplified. For instance, it is expected [123] that the setup realizing the term $\sim\cos\left(3\phi\right)$ [Fig. 8(b)] has enough free parameters to implement the potential without the additional inclusion of a $\cos(\phi)$ and a $\cos(2\phi)$ junction.
[…]
FIG. 8: Realization of a junction with $\cos(2\phi)$ energy profile (transporting pairs of Cooper pairs, see also Ref. [101]) in panel a). A $\cos(3\phi)$ element is realized according to the layout in b). In the circuit diagrams above for the generalized Josephson junctions, the Josephson junctions have the same parameters and so do the inductors. The shunt capacitance ($\epsilon_{C}$) is the dominant energy scale.
At any rate, the above potential has the sought-after minima with different curvatures. This circuit element is therefore at low energies described as a type of quantum inductor, which can exhibit superpositions of different inductance values. Using such an element to connect nodes $j$ and $j-1$, this element can be described by the energy
$(E_{L}+\delta E_{L}\sigma_{j}^{z})(\phi_{j}-\phi_{j-1})^{2}\ ,$ (28)
where the relative difference of the inductances is given by $\delta E_{L}/E_{L}=2\delta$ [with $\delta$ as it appears in Eq. (27)], and the pseudo-spin $\sigma^{z}$ represents the two distinct quantum well states of the inductor, $\left|0_{j}\right>$ and $\left|1_{j}\right>$, having eigenvalue $-1$ and $+1$, respectively. We note that this description is of course a crude approximation. In particular, finite capacitances will likely induce transitions $\sim\sigma^{x}_{j}$. While we for now discard this possibility for simplicity, we will outline in the outlook, that the inclusion of coherent spin-flip processes (or even nonlocal spin-spin interactions) by no means pose a danger to the proposal outlined here, but to the contrary, render it even more interesting (as they may give rise to a rudimentary simulation of an analog quantum spacetime with an intrinsic coherent dynamics of its own).
The full proposal to realize and measure superpositions of spacetime geometries now works as follows. Consider the circuit in Fig. 9 consisting of three regions, two regions where the superconducting nodes are connected by Josephson junctions and a third one sandwiched between the the other two where superconducting nodes are connected by the newly introduced quantum inductors. This system is represented by the following Hamiltonian
$\tilde{\mathcal{H}}=$ $\sum_{j}E_{C}N_{j}^{2}+\sum_{j=j^{\prime}}^{j^{\prime\prime}}\left(E_{L}+\delta E_{L}\sigma_{j}^{z}\right)\left(\phi_{j+1}-\phi_{j}\right)^{2}$
$+\sum_{j\notin[j^{\prime},j^{\prime\prime}]}E_{L}^{\prime}\left(\phi_{j+1}-\phi_{j}\right)^{2}\ .$ (29)
For our proposal the circuit needs to be prepared in the maximally entangled state $\left(\left|\tilde{0}\right>+\left|\tilde{1}\right>\right)/\sqrt{2}$, where
$\left|\tilde{0}\right>$ $=\left|0_{j^{\prime}}0_{j^{\prime}+1}...0_{j^{\prime\prime}-1}0_{j^{\prime\prime}}\right>\equiv\left(\begin{array}[]{c}1\\ 0\end{array}\right)$
$\left|\tilde{1}\right>$ $=\left|1_{j^{\prime}}1_{j^{\prime}+1}...1_{j^{\prime\prime}-1}1_{j^{\prime\prime}}\right>\equiv\left(\begin{array}[]{c}0\\ 1\end{array}\right).$
To measure this entanglement we will exploit the fact that the two states of the asymmetric potential correspond to different signal speeds, present in the sandwiched region. Let us now initialize a wave packet in the trivial region to the left. Assuming for concreteness a signal with large amplitude in $\phi$, we can simply describe it with a wave function $\phi(x-ut)$ (using a continuous space representation $x=\Delta xj$, valid if the wave packet is large with respect to the lattice size $\Delta x$), centered around a position on the left, and moving towards the region with the quantum inductors. The propagation speed here is $u=\Delta x\sqrt{E_{C}E_{L}^{\prime}}$. Within this region, the signal will now propagate with two different speeds. If the quantum inductor region is large enough with respect to the speed difference and the wave packet size, we will create a fully entangled state. The transition from the incoming wave packed to the left of the quantum inductor region, and the outgoing wave packet to the right of the quantum inductor region can be described as
$\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}\phi\left(x-ut\right)\\ \phi\left(x-ut\right)\end{array}\right)\rightarrow\frac{1}{\sqrt{2}}\left(\begin{array}[]{c}\phi\left(x+x_{0}-ut\right)\\ \phi\left(x+x_{1}-ut\right)\end{array}\right)\ .$ (34)
The positions $x_{0}$ and $x_{1}>x_{0}$ (i.e., the wave packet ${\phi(x+x_{1}-ut)}$ is the further advanced one) express the different positions of the outgoing wave packets due to the different times of flight within the quantum inductor region. If the overlap vanishes, i.e. if ${\phi(x+x_{0}-ut)\phi(x+x_{1}-ut)\approx 0}$, the entanglement is maximal.
[…]
FIG. 9: Illustration of quantum superposition of spacetime geometry and its detection with travelling signals in snapshots. a) Initializing a signal on the left hand side (with ordinary classical flat spacetime), the signal starts to separate in the quantum spacetime region due to different time of flight, and ultimately exits in the normal region on the right in a fully entangled state. b) After pplying a Hadamard gate on the maximally entangled state in the quantum spacetime region, and letting the signals reflect at a hard wall (yellow stars indicate the reflection event) the signals interfere either constructively or descructively.
A correlated measurement of the wave packet (e.g., time of flight measurement) and the state of the circuit $\left(\left|\tilde{0}\right>\text{ or }\left|\tilde{1}\right>\right)$ provides information about the presence of the entanglement. Note that in order to verify the quantum coherent nature of the analog spacetime, one can easily formulate a more sophisticated interference setup. To this end, after the wave packets have traversed the quantum inductor region, we perform a Hadamard gate on the qubit states, and add a hard wall on the far right end, such that the further advanced wave packet can reflect. This sequence yields the state
[…]
where in the last line, the updated shift $\widetilde{x}_{1}$ takes into account the distance travelled up to and after collision with the hard wall. As can be seen on the last line of the above equation, the further advanced package now travels backwards ($+ut$ instead of $-ut$), and is thus bound to interfere with the delayed wave packet (which still travels forward, $-ut$). In this setup, interference is constructive for the $\left|\tilde{0}\right>$ branch, and destructive for the $\left|\tilde{1}\right>$ branch. Hence, by projective measurement of the signal either directly after passing the quantum inductor region, or measuring after the succession of Hadamard gate and reflection, the quantum nature of the spacetime geometry can be unequivocally probed.
VIII Conclusions and outlook
@@
In this paper we have proposed using quantum circuits to implement two aspects of quantum gravity: 1) analog horizons and 2) quantum superpositions of spacetime geometries. Even though the former idea is a well explored concept in solid state systems in general, we here explicitly demonstrate the capacity of superconducting circuits to emulate arbitrary spacetime configurations. By identifying a minimal fundamental set of necessary circuit elements, we further unravel a number of surprising findings pertinent to lattice systems, which allow in particular to create horizons with trans-Planckian curvature, allowing us to disentangle the effect of curvature of background metric from the effect of a horizon – and in doing so, exploring the extreme regime of horizons with diverging curvatures, which as far as we know has not been done in any existing proposal. To implement the analog horizons in a quantum circuit, we proposed a general way of creating negative inductances with Josephson junctions transiently driven by nearby current loops. We further uncovered two subtly distinct ways to engineer a region with overtilted dispersion relation, either with inductive coupling of nearest or next-to-nearest neighbour nodes. While for the latter, horizons have a combination of black and white hole character (as is the most usual case in lattices), the former [Eq.(III)] hosts two distinct kinds of horizons, one that acts purely as a black hole horizon and the other as a white hole horizon. For this second type of circuit, it is not only the horizons, but also the region between the horizons (wormhole interior) that is unstable due to the presence of an exceptional point. This instability also contributes to the observed radiation (in addition to the evaporation of the horizons) as can be confirmed by tracking the quantum fluctuations of phase difference and conjugate charge with time (Fig. 5). In addition to Bogoliubov diagonalization, we used an extension of Klich’s trace-determinant formula to obtain the correlation functions numerically. The proof of this extension remains elusive for a general operator, but in the parameter regime where the generalised Bogoliubov transformation is applicable, the results of both numerical methods agree.
Our second proposal concerns the controversial idea of superpositions of spacetime geometries. This idea was shown to emerge as a nontrivial extension to the above negative inductance realization. The emulation of such geometries relies on engineering junction elements with an energy-phase relationship that exhibits an asymmetric double well potential (Eq.(27)), so-called multi-stable Josephson junctions. The different curvatures of the two minima of the potential give us the two inductances that act as the basis states for the quantum inductance. We proposed a setup where a classical signal traverses the quantum spacetime region, resulting in a superposition of two different signal speeds. The superposition can be measured by making correlated measurements of the quantum inductance and the signal moving through the circuit, or by means of a delayed choice experiment.
@@
In this paper, in addition to standard circuit elements, we have also included elements that are still being developed experimentally (such as quantum gyrators), or introduced new effective circuit elements (based on regular or multistable Josephson junctions), which, even though experimentally tested in existing works, are here proposed to be operated in previously unexplored regimes. We here therefore point at feasible small scale experimental tests that can act as proofs of principle for the basic physical principles behind the considered physics. For instance, our proposal of using current loops to invert the effective inductance of a Josephson junction can be tested on a single transmon, where the qubit frequency becomes imaginary $\sqrt{E_{J}E_{C}}\to\sqrt{-E_{J}E_{C}}$ for transient times. The resulting accumulation of phase or charge quantum fluctuations (or limits thereof due to coupling to the environment) can thus be tested in an “imaginary” qubit device. Such a study might be also of some fundamental interest, when extending the description of the junction beyond the quadratic approximation, where the expansion around the energy minimum or the energy maximum can be regarded as real and imaginary twins of nonlinear dynamics – ready to be explored in future works. For the case of quantum inductors, one can likewise consider a simple circuit with either one or a few quantum inductors in series, where the low energy excitations induced in the two minima will have different oscillation frequency – allowing for the engineering of (nearly) harmonic oscillators in a quantum superposition of two resonance frequencies.
Finally, we comment on extensions towards a more fully-fletched simulation of analog quantum spacetimes. Namely, in our treatment in the main text, the states of the quantum inductors had, as of now, no intrinsic dynamics. While this is the simplest possilbe treatment from a theoretical stand point, in experiment it would require a tuning to very extreme ratios of small charging energy compared to the junction energy scale. For more moderate ratios, it cannot be excluded that the basis states of the quantum inductors receive an intrinsic dynamics, for example a tunnel coupling between the two minima, $\sigma_{j}^{x}$, or non-local couplings such as $\sigma_{j}^{\alpha}\sigma_{j^{\prime}}^{\beta}$ ($\alpha,\beta=x,y,z$), which have to be added in the Hamiltonian in Eq.(29). In this regard, note that this specific quantum spacetime proposal is in line with currently existing plans to engineer quantum computing hardware by means of flux qubits based on multi-stable Josephson junctions [100, 101, 102]. Crucially, however, while the above described type of crosstalk is obviously detrimental for the performance of the hardware when used for standard quantum information applications, for the here considered analog quantum gravity simulation, such imperfections are by no means a disadvantage – to the contrary, they may open up a new road to explore analog quantum gravity where the spacetime metric has its quantum coherent dynamics.
IX Acknowledgements
@@
We acknowledge interesting and fruitful discussions with Zaki Leghtas, Roudy Hanna, Philippe Campagne-Ibarcq and David DiVincenzo. This work has been funded by the German Federal Ministry of Education and Research within the funding program Photonic Research Germany under the contract number 13N14891.

### v4

@@ 6 Perspective on wormhole evaporation over long times
As shown in the previous section, charge and phase fluctuations accumulate very rapidly due to the fundamental instability inherent to the wormhole horizons or interior (in case of nearest neighbour inductive coupling) – an effect distinct from regular Hawking radiation. In particular, the more the system accumulates charge and phase fluctuations over time, the more the quadratic approximation of the Josephson energy becomes inaccurate, such that predicting the system dynamics for long times becomes a much harder task (as the system can no longer be approximated by non-interacting bosons). This is well beyond the scope of the present work. We nonetheless find it illustrative to speculate on a qualitative level about the long term fate of the wormhole – especially as it allows us to distinguish between intrinsic (spontaneous) collapse, and dissipative relaxation. In addition, some of the resulting concepts provide an interesting segue to the subsequent idea: quantum superpositions of the spacetime geometry.
@@
To this end, let us take into account interactions with an environment. For the sake of simplicity and concreteness, we consider the phase difference $\phi_{j}-\phi_{j-1}+\phi_{\text{ext},j}$ across a single junction. At the beginning of the wormhole quench, this phase difference is either localized around 0 (minimum of the cosine, if the junction is positioned outside the wormhole) or around $\pm\pi$ (maximum of the cosine, inside the wormhole). In the presence of the instability, the quantum fluctuations of this phase difference blow up over time, as shown in the previous section. Now suppose that the environment entangles (either weakly or strongly) with the current across this junction. This means that the phase difference gets spontaneously projected onto a more localized state. Notably, this type of environment-induced process thus diminishes quantum fluctuations. We note that the same picture holds in perfect analogy for the accumulation of charge noise quantum fluctuations, respectively, the reduction thereof by a dissipative process extracting information about the charge.
@@
To proceed, note that the location of the wave function now is no longer at 0 or $\pi$, but can be (with a finite probability) at a certain distance from the minimum or maximum. With the new location updated, we again quadratically expand the cosine around the new peak position of the phase difference. Importantly, due to the cosine behaviour of the junction the resulting effective inductance is now different (the curvature of the cosine obviously changes as a function of the position at which it is calculated). Thus, as time progresses, and the phase difference starts to classically diffuse after repeated environment-induced measurements, the effective inductance at a given junction likewise fluctuates classically. At this point, the impact of such dissipative processes could likewise be detectable in time-of-flight measurements of wave packets within the array. At any rate, the most likely long-term outcome is that all phase differences relax to the ground state, equivalent to the complete evaporation of the analog wormhole. We note that for a system on a ring (with periodic boundaries), or a system with a finite mass term (Cooper-pair leakage to ground), the emergence of soliton states might be a possibility.
@@
Let us repeat that even though the environment will undoubtedly have a big impact on the system dynamics (especially beyond the immediate transient time scale), the experimental signatures of quantum collapse (due to complex eigenvalues) and classical collapse (due to dissipation) could not be more different. The instability due to the inhomogeneous metric leads to an increase of quantum fluctuations, whereas the environment induces classical fluctuations. Thus, the observation of the former can distinguish between how the system intrinsically reacts to the creation of an instability, as compared to the impact of external perturbations.
@@
To conclude this section, we consider the very hypothetical case of negligible coupling to the environment, as it reveals an interesting additional idea which we hope to pursue in future works. If we assume that the build-up of quantum fluctuations of the phase could progress unimpeded over longer times, the nonlinearity of the Josephson energies results not in classical fluctuations of the effective inductance (as in the previous paragraph), but in a system that can be interpreted as being in a quantum superposition of different effective inductances – a form of quantum superposition of the spacetime geometry itself. However, within the above considered setting, this effect will likely not survive for very long times due to environment-induced decoherence of the quantum fluctuations, and even for a highly protected system, it would be hard to analyse the effect in an unambiguous way. In upcoming papers, we plan to provide a modification of this idea, not relying on instabilities (and thus not requiring any ultra-fast quenches), but instead on more general notions of Josephson junctions with multiple minimas, so-called multi-stable Josephson junctions [118, 119, 120]. This endeavour will on the one hand introduce the notion of a quantum inductance, and with it, a likely more stable version of a quantum superposition of the analog spacetime. As we will show, there exist protocols allowing classical signals to entangle with quantum spacetime, marking a highly non-trivial form of backaction between fields and the spacetime they traverse.
7 Conclusions and outlook
@@
In this paper we have proposed using quantum circuits to implement analog event horizons on a discrete lattice. Even though this idea is a well explored concept in solid state systems in general, we here explicitly demonstrate the capacity of superconducting circuits to emulate arbitrary spacetime configurations. By identifying a minimal fundamental set of necessary circuit elements, we further unravel a number of surprising findings pertinent to lattice systems, which allow in particular to create horizons where the change in metric parameters happens at a level of a single lattice site. To implement the analog horizons in a quantum circuit, we proposed a general way of creating negative inductances with Josephson junctions transiently driven by nearby current loops. We further uncovered two subtly distinct ways to engineer a region with overtilted dispersion relation, either with inductive coupling of nearest or next-to-nearest neighbour nodes. While for the latter, horizons have a grey hole character (as is the most usual case in lattices), the former [Eq. (15)] hosts two distinct kinds of horizons, one that acts purely as a black hole horizon and the other as a white hole horizon. For this second type of circuit, it is not only the horizons, but also the region between the horizons (wormhole interior) that is unstable due to the presence of an EP. This instability also contributes to the observed radiation (in addition to the evaporation of the horizons) as can be confirmed by tracking the quantum fluctuations of phase difference and conjugate charge with time (Fig. 6). In addition to Bogoliubov diagonalization, we used an extension of Klich’s trace-determinant formula to obtain the correlation functions numerically. The proof of this extension remains elusive for a general operator, but in the parameter regime where the generalised Bogoliubov transformation is applicable, the results of both numerical methods agree.
@@
In this paper, in addition to standard circuit elements, we have also included elements that are still being developed experimentally (such as quantum gyrators), or introduced new effective circuit elements (negative inductors based on regular Josephson junctions), which, even though experimentally tested in existing works, are here proposed to be operated in previously unexplored regimes. We here therefore point at feasible small scale experimental tests that can act as proofs of principle for the basic physical principles behind the considered physics. For instance, our proposal of using current loops to invert the effective inductance of a Josephson junction can be tested on a single transmon, where the qubit frequency becomes imaginary $\sqrt{E_{J}E_{C}}\to\sqrt{-E_{J}E_{C}}$ for transient times. The resulting accumulation of phase or charge quantum fluctuations (or limits thereof due to coupling to the environment) can thus be tested in an “imaginary” qubit device. Such a study might be also of some fundamental interest, when extending the description of the junction beyond the quadratic approximation, where the expansion around the energy minimum or the energy maximum can be regarded as real and imaginary twins of nonlinear dynamics – ready to be explored in future works.
Acknowledgements
@@
We acknowledge interesting and fruitful discussions with Zaki Leghtas, Roudy Hanna, Philippe Campagne-Ibarcq and David DiVincenzo.
Funding information
This work has been funded by the German Federal Ministry of Education and Research within the funding program Photonic Research Germany under the contract number 13N14891.